OpenStudy (ashking1):

Write the sum or difference in the standard form a + bi. ( 3 + 3i) - ( -6 + i)

OpenStudy (solomonzelman):

first, distribute the - next to parenthesis

OpenStudy (solomonzelman):

- ( -6 + i) = ?

OpenStudy (ashking1):

i did that already

OpenStudy (solomonzelman):

what did you get/

OpenStudy (solomonzelman):

?

OpenStudy (ashking1):

(3+3i)(6-i)

OpenStudy (solomonzelman):

- ( -6 + i) = + 6 - i (because you multiply every term in the parenthesis times -1)

OpenStudy (solomonzelman):

So, ( 3 + 3i) - ( -6 + i) = ( 3 + 3i) + 6 - i = 3 + 3i + 6 - i = 3 + 6 + 3i - i = ?? (you tell me)

OpenStudy (ashking1):

9+ 4i but these are my choices and it does not match -9 - 2i -3 + 4i 9 - 2i 9 + 2i

OpenStudy (solomonzelman):

yes it does match, you just need to finish what I have started

OpenStudy (solomonzelman):

3+6=? 3i-i=? so 3+6+3i-i=?

OpenStudy (ashking1):

oh ok 9+2i

OpenStudy (ashking1):

thank you can you help me with one more